# MMSI-Bench: A Benchmark for Multi-Image Spatial Intelligence

### v1

@@ B.5 License and Access
Appendix C Comparison with Other Spatial Question-Answering Benchmarks
@@
VSR [34] Auto & Human Single-Image 10K Indoor & Outdoor 25.0%
BLINK [22] Auto & Human Single/Multi-Image 3,807 (572) Real-world & Synthetic 44.4%
UniQA-3D [68] Auto & Human Single/Multi-Image 500 Real-world & Synthetic 14.2%
MuriBench [57] Auto & Human Multi-Image 2,600 (478) Real-world & Synthetic 25.2%
SpatialVLM [15] Auto & Human Single-Image 546 Indoor & Outdoor <30%
SpatialRGPT [17] Auto Single-Image 1,406 Real-world & Synthetic <42%
MMIU [40] Auto Multi-Image 11K (2904) Real-world & Synthetic <45%
LEGO-Puzzle [55] Human Single/Multi-Image 1K Synthetic 39.5%
VSI-Bench [65] Auto & Human Video 5K Indoor 33%
MMSI-Bench Human Multi-Image 1,000 Indoor & Outdoor 56.2%
Table 3: Comparison of spatial question-answering benchmarks. Some benchmarks contain QA pairs outside spatial understanding; the number of spatial QA is shown in parentheses. Human-AI performance gap is defined as the accuracy difference between the best-performing model reported in the corresponding paper and human performance; for benchmarks without reported human results, the gap is calculated with respect to 100% accuracy and is marked as “<<<value%”. Note: Scenario entries labeled as “Indoor & Outdoor” or “Indoor” refer to real-world images.
Despite significant advances in multimodal large language models (MLLMs), robust evaluation of their spatial intelligence—especially in multi-image, real-world scenarios—remains limited. As summarized in Table 3, most existing benchmarks focus narrowly on single-image spatial understanding or offer only limited multi-image coverage as minor components within more general VQA suites (e.g., BLINK [22], MuriBench [57]), rather than serving as dedicated spatial reasoning benchmarks. Furthermore, several datasets rely heavily on template- or rule-based question generation [50, 68, 40, 65], which restricts both the diversity and depth of spatial reasoning challenges. Synthetic settings such as LEGO-Puzzle [55] further lack the complexity and contextual richness of real-world environments.
@@
Appendix D Additional Implementation Details and Experimental Results
D.1 Implementation Details for Model Evaluation
@@
Table 4: Prompts used for different evaluation settings.
For the visual prompting strategy, “{image with visual prompt}” refers to the image input combined with a visual prompt drawn using PATS [42], as illustrated in Figs. 891011.
@@
The prompt used for evaluating reasoning processes with a human-annotated reference is shown in Table 6.
@@
{human-annotated reference reasoning process} You are given images, a question about the images, the correct answer, a human-annotated reference reasoning process, and a reasoning process to be evaluated. Your task is to evaluate the reasoning process and determine whether it is entirely correct. If there are any errors, identify the type of error that most critically led to the incorrect answer by selecting the corresponding category number. If the reasoning process is completely correct, select only category 5. Categories:
@@
Table 6: Prompt used for evaluating reasoning processes with human-annotated reference in our study.
D.2 Human Evaluation Setup
@@
D.3 Complete Evaluation Results of Prompting Techniques on MMSI-Bench
@@
Appendix E Details of the Data Curation Process
E.1 Dataset Sources
@@
ScanNet [18] : ScanNet is a large-scale RGB-D video dataset designed for 3D scene understanding in real-world indoor environments. It consists of 2.5 million RGB-D images across 1,513 scans from 707 unique spaces, each annotated with 3D camera poses, surface reconstructions, and dense object-level semantic segmentations. The dataset was collected using a scalable pipeline that combines easy data capture with crowdsourced annotations, making it one of the most richly-labeled and comprehensive resources for training and evaluating deep learning models on tasks such as 3D object classification and semantic voxel labeling.
@@
nuScenes [13]: nuScenes is a multimodal autonomous driving dataset that features a complete sensor suite, including 6 cameras, 5 radars, and 1 lidar, all with 360-degree coverage. It contains 1,000 driving scenes, each 20 seconds long, fully annotated with 3D bounding boxes for 23 object classes and 8 attributes. nuScenes stands out for its diversity of sensor data, large scale of annotations, and support for advanced tasks such as 3D detection and tracking, making it a benchmark for research in perception for autonomous vehicles.
@@
Matterport3D [14]: Matterport3D is a large-scale RGB-D dataset capturing 10,800 panoramic viewpoints from 194,400 images of 90 building-scale indoor scenes. It provides surface reconstructions, globally aligned camera poses, and detailed 2D and 3D semantic segmentations at both region and object levels. The comprehensive panoramic coverage and precise global alignment enable a wide range of computer vision tasks, including keypoint matching, semantic segmentation, and view overlap prediction, particularly in the context of indoor scene understanding.
@@
Ego4D [23]: Ego4D is an unprecedented egocentric video dataset, comprising 3,670 hours of first-person footage collected by 931 participants across 74 locations in 9 countries. The dataset covers a wide array of daily-life scenarios and includes supplementary data such as audio, 3D meshes, gaze, and multi-camera views. With millions of annotations supporting various benchmark tasks, Ego4D enables research into first-person perception, episodic memory, and social interactions, significantly advancing the field of egocentric visual understanding.
@@
AgiBot-World [9]: AgiBot-World is a large-scale robotic manipulation dataset featuring over 1 million trajectories across 217 tasks in five real-world domains—including domestic, retail, industrial, restaurant, and office environments—collected by more than 100 real robots. Data collection is standardized and includes human-in-the-loop verification to ensure high quality and diversity. The dataset supports a wide range of hardware, from simple grippers to dexterous hands with visuo-tactile sensors, and provides rich multimodal annotations. AgiBot-World enables the development and evaluation of generalist robot policies in complex, long-horizon, and diverse scenarios, moving towards scalable, general-purpose robotic intelligence.
@@
DTU [26]: The DTU dataset is a large-scale multi-view stereo (MVS) benchmark designed for 3D reconstruction research. It contains 80 scenes with significant variability in surface properties and geometric complexity. Each scene is captured from 49 or 64 accurate camera positions and includes high-quality reference scans obtained via structured light. This dataset advances the scale and diversity compared to previous MVS datasets and serves as a challenging benchmark for evaluating the accuracy and completeness of stereo reconstruction algorithms.
@@
DAVIS 2017 [48]: DAVIS 2017 is a benchmark dataset for video object segmentation, consisting of 150 video sequences and a total of 10,459 annotated frames with pixel-level masks for multiple objects per scene. Compared to earlier versions, DAVIS 2017 introduces greater complexity, including multiple objects, occlusions, small and fine structures, and fast motion. The dataset has driven significant progress in video object segmentation and is widely used for benchmarking and competition in the field.
@@
Waymo Open Dataset [54]: The Waymo Open Dataset is one of the largest and most diverse multimodal datasets for autonomous driving. It comprises 1,150 scenes, each lasting 20 seconds, captured across multiple cities with extensive geographic coverage. Each scene includes synchronized and calibrated high-resolution camera images and LiDAR data, along with exhaustive 2D and 3D bounding box annotations and consistent object tracking IDs. The dataset supports research in detection, tracking, and sensor fusion, providing a valuable foundation for developing robust and generalizable self-driving technologies.
The distribution of data sources used for each question in MMSI-Bench is shown in Table 8.
Category Count Percentage
Matterport3D 463 46.30%
ScanNet 280 28.00%
Ego4D 67 6.70%
AgiBot-World 45 4.50%
DTU 72 7.20%
nuScenes 39 3.90%
DAVIS 2017 29 2.90%
Waymo 5 0.50%
Table 8: Sample counts and proportions for each category
E.2 Data Annotation and Quality Control Interfaces
@@ E.3 General Guidelines
E.4 Data Format and Structure
@@
E.5 Quality Control and Validation
@@
Appendix F Representative MMSI-Bench Samples from Each Category (Full Version)

### v2

@@ B.6 License and Access
Appendix C Motivation for Multi-Image Input
We expect that general spatial intelligence will, like humans, operate on video inputs. However, we start with a multi-image setting for three reasons:
Foundational to video. Multi-image understanding is the basis of video understanding. Analogous to SLAM, a system reconstructs a scene and estimates camera pose over a stream, but the front-end module (VO) still estimates the delta pose between consecutive frames before constructing a holistic perception of the scene. Given current limitations of MLLMs in spatial understanding, models must first handle multiple images before they can reliably understand full videos.
Clearer diagnosis. Multiple images allow more systematic analysis of current MLLMs’ failure modes. With videos, if a model answers incorrectly, it is hard to localize the cause, e.g., long-video limits, image retrieval over long sequences, or spatial reasoning itself. The multi-image setup is more spatially focused and easier for error analysis, which we believe accelerates community progress.
Relevant proxy for video. Multi-image benchmarks also reflect capabilities needed for video-based systems. Our hypotheses are: (i) although humans learn spatial intelligence from videos, they can solve MMSI-Bench, suggesting our tasks capture a necessary condition for general (video-based) spatial intelligence; and (ii) a multi-image setting can be viewed as a multi-frame, low-FPS special case of video.
We hope MMSI-Bench will inspire future video-based benchmarks.
Appendix D Comparison with Other Spatial Question-Answering Benchmarks
Table 4: Comparison of spatial question-answering benchmarks. Some benchmarks contain QA pairs outside spatial understanding; the number of spatial QA is shown in parentheses. Human-AI performance gap is defined as the accuracy difference between the best-performing model reported in the corresponding paper and human performance; for benchmarks without reported human results, the gap is calculated with respect to 100% accuracy and is marked as “<<value%”. Note: Scenario entries labeled as “Indoor & Outdoor” or “Indoor” refer to real-world images.
@@ Appendix D Comparison with Other Spatial Question-Answering Benchmarks
VSR (Liu et al., 2023) Auto & Human Single-Image 10K Indoor & Outdoor 25.0%
BLINK (Fu et al., 2024) Auto & Human Single/Multi-Image 3,807 (572) Real-world & Synthetic 44.4%
UniQA-3D (Zuo et al., 2024) Auto & Human Single/Multi-Image 500 Real-world & Synthetic 14.2%
MuriBench (Wang et al., 2024a) Auto & Human Multi-Image 2,600 (478) Real-world & Synthetic 25.2%
SpatialVLM (Chen et al., 2024a) Auto & Human Single-Image 546 Indoor & Outdoor <30%
SpatialRGPT (Cheng et al., 2024) Auto Single-Image 1,406 Real-world & Synthetic <42%
MMIU (Meng et al., 2024) Auto Multi-Image 11K (2904) Real-world & Synthetic <45%
ERQA (Team et al., 2025b) Human Single/Multi-Image 400 Indoor & Outdoor <35%
LEGO-Puzzle (Tang et al., 2025) Human Single/Multi-Image 1K Synthetic 39.5%
VSI-Bench (Yang et al., 2024) Auto & Human Video 5K Indoor 33%
MMSI-Bench Human Multi-Image 1,000 Indoor & Outdoor 55.3%
Despite significant advances in MLLMs, robust evaluation of their spatial intelligence—especially in multi-image, real-world scenarios—remains limited. As summarized in Table 4, most existing benchmarks focus narrowly on single-image spatial understanding or offer only limited multi-image coverage as minor components within more general VQA suites (e.g., BLINK (Fu et al., 2024), MuriBench (Wang et al., 2024a)), rather than serving as dedicated spatial reasoning benchmarks. Furthermore, several datasets rely heavily on template- or rule-based question generation (Ray et al., 2024, Zuo et al., 2024, Meng et al., 2024, Yang et al., 2024), which restricts both the diversity and depth of spatial reasoning challenges. Synthetic settings such as LEGO-Puzzle (Tang et al., 2025) further lack the complexity and contextual richness of real-world environments.
@@
To further quantify the linguistic diversity of MMSI-Bench, we conduct a comparative analysis against representative non-single-image spatial reasoning benchmarks, including VSI-Bench, the multi-image subset of ERQA, and LEGO-Puzzle. We analyze the complexity of the questions across these datasets at the vocabulary, phrase, and syntactic levels. The results are summarized in Table 5.
The analysis demonstrates that MMSI-Bench exhibits superior diversity across all metrics. At the vocabulary level, MMSI-Bench features 1,108 distinct words, significantly more than the other benchmarks. At the phrase level, it leads in the number of unique 3-grams and 5-grams, suggesting a wider variety of phrasal combinations and expressions. Most notably, at the syntactic level, MMSI-Bench contains 993 unique syntactic structures—nearly double that of VSI-Bench and more than triple that of ERQA. This linguistic and structural complexity stems from our fully human-curated annotation process, which is driven by experts and avoids any reliance on templates. This ensures that models are challenged not only with spatial reasoning tasks but also with understanding and processing diverse, naturally-phrased questions that reflect real-world linguistic distribution. Consequently, it sets a higher bar for evaluating a model’s generalization capabilities.
Table 5: Quantitative comparison of question diversity against major non-single-image spatial reasoning benchmarks. Distinct words: Counts the total number of unique words after lowercasing and tokenization across all questions, measuring vocabulary richness. Unique n-grams: Counts the number of unique consecutive n-word sequences (here, n=3 and 5) after tokenization, measuring phrasal diversity. Syntactic structures: Calculated by parsing the syntax of each sentence, abstracting it into a tree structure by removing specific words, and then counting the number of unique structures. This measures the diversity of sentence constructions.
Statistic VSI-Bench MMSI-Bench ERQA (Multi-Img) LEGO-Puzzle
Distinct words 398 1,108 559 275
Unique 3-grams 4,482 8,499 2,483 1,023
Unique 5-grams 12,972 13,010 3,318 1,373
Syntactic structures 473 993 310 51
Appendix E Additional Implementation Details and Experimental Results
E.1 Benchmark models.
We conduct a comprehensive evaluation of a wide range of MLLMs, encompassing both proprietary and open-source families, diverse model scales, and recent architectural advances. For proprietary models, we includes GPT-5 (OpenAI, 2025c), o3 (OpenAI, 2025d), GPT-4.5 (OpenAI, 2025b), GPT-4.1 (OpenAI, 2025a), GPT-4o (OpenAI, 2024), Gemini-2.5-Pro (DeepMind, 2025), Claude-3.7-Sonnet (Anthropic, 2025), dots.vlm 1 (Huo et al., 2025), GLM-4.5V-thinking (Hong et al., 2025), Doubao-1.5-pro (Seed, 2025), and Seed1.5-VL (Guo et al., 2025). On the open-source side, we systematically assess state-of-the-art models such as Qwen2.5-VL (Bai et al., 2025), LLaVA-OneVision (Li et al., 2024), InternVL3 (Zhu et al., 2025), InternVL2.5 (Chen et al., 2024b), DeepSeek-VL2 (Wu et al., 2024), Llama-3.2-11B-Vision (Meta, 2024), and NVILA (Liu et al., 2024c).
E.2 Implementation Details for Model Evaluation
We report accuracy (%) using exact match between answers extracted from model outputs and ground-truth answers for our multiple-choice questions. If a model fails to generate an answer in the required format, we adopt the LLM-based fallback strategy from VLMEvalKit (Duan et al., 2024) to extract the intended response.
@@ E.2 Implementation Details for Model Evaluation
Table 6: Prompts used for different evaluation settings.
@@
For the visual prompting strategy, “{image with visual prompt}” refers to the image input combined with a visual prompt drawn using PATS (Ni et al., 2023), as illustrated in Figs. 891011.
@@
The prompt for evaluating reasoning processes with a human-annotated reference is shown in Table 8.
Table 8: Prompt used for evaluating reasoning processes with human-annotated reference in our study. Here, the structured format of the reasoning process means that, based on the original human-annotated reasoning steps, each step is further labeled with its reasoning type. There are four reasoning types, Grounding, Overlap-Matching and Scene-Reconstruction, Situation-Transformation Reasoning, and Spatial-Logic Reasoning, which correspond to four types of possible errors, respectively. For example, a reasoning process can be transformed into the structured format as follows: “The pink bench is to the east of the bed. The white wall lies to the south. Thus, the bathroom lies in the southeast corner of the sleeping region.” will become “The pink bench is to the east of the bed (Grounding). The white wall lies to the south (Grounding). Thus, the bathroom lies in the southeast corner of the sleeping region (Overlap-Matching and Scene-Reconstruction).”
@@
{human-annotated reference reasoning process with structured format} You are given images, a question about the images, the correct answer, a human-annotated reference reasoning process, and a reasoning process to be evaluated. Your task is to evaluate the reasoning process and determine whether it is entirely correct. If there are any errors, identify the type of error that most critically led to the incorrect answer by selecting the corresponding category number. If the reasoning process is completely correct, select only category 5. Categories:
@@
E.3 Human Evaluation Setup
@@
E.4 Fine-tuned Spatial-Reasoning Models vs. Base Models
We compare recent spatially fine-tuned models with their base versions on MMSI-Bench (Table 9). Overall, gains are small and inconsistent across tasks; in several cases, fine-tuned models match or underperform the base models. A likely reason is that their training data, built from limited templates, fails to generalize to the diverse and challenging scenarios in MMSI-Bench, pointing to the need for richer data or improved training/architectures.
Table 9: Performance of models fine-tuned on spatial-reasoning data and their base models on MMSI-Bench.
Models Positional Relationship Attribute Motion MSR Avg.
Cam.–Cam. Obj.–Obj. Reg.–Reg. Cam.–Obj. Obj.–Reg. Cam.–Reg. Meas. Appr. Cam. Obj.
Qwen2.5-VL 3B 26.9 27.7 30.9 29.1 28.2 34.9 31.2 16.7 17.6 27.6 23.2 26.5
Spatial-MLLM 3B (Wu et al., 2025a) 23.6 26.6 37.0 34.9 32.9 24.1 18.8 34.9 34.9 25.0 28.8 27.7
InternVL2_5-8B 32.3 27.7 29.6 32.6 24.7 32.5 26.6 27.3 16.2 31.6 30.3 28.7
InternSpatial (Deng et al., 2025) 20.4 27.6 28.4 29.0 30.5 26.5 29.6 27.2 14.8 23.6 31.3 26.9
Qwen2.5-VL-32B 24.7 26.6 29.6 22.1 32.9 31.3 31.2 24.2 18.9 35.5 27.8 27.7
RoboBrain2.0 32B (Team et al., 2025a) 21.5 28.7 27.1 22.0 27.0 32.5 31.2 36.3 36.4 23.6 29.2 28.5
E.5 Complete Evaluation Results of Prompting Techniques on MMSI-Bench
@@
Appendix F Details of the Data Curation Process
F.1 Dataset Sources
@@
ScanNet (Dai et al., 2017): ScanNet is a large-scale RGB-D video dataset designed for 3D scene understanding in real-world indoor environments. It consists of 2.5 million RGB-D images across 1,513 scans from 707 unique spaces, each annotated with 3D camera poses, surface reconstructions, and dense object-level semantic segmentations. The dataset was collected using a scalable pipeline that combines easy data capture with crowdsourced annotations, making it one of the most richly-labeled and comprehensive resources for training and evaluating deep learning models on tasks such as 3D object classification and semantic voxel labeling.
@@
nuScenes (Caesar et al., 2020): nuScenes is a multimodal autonomous driving dataset that features a complete sensor suite, including 6 cameras, 5 radars, and 1 lidar, all with 360-degree coverage. It contains 1,000 driving scenes, each 20 seconds long, fully annotated with 3D bounding boxes for 23 object classes and 8 attributes. nuScenes stands out for its diversity of sensor data, large scale of annotations, and support for advanced tasks such as 3D detection and tracking, making it a benchmark for research in perception for autonomous vehicles.
@@
Matterport3D (Chang et al., 2017): Matterport3D is a large-scale RGB-D dataset capturing 10,800 panoramic viewpoints from 194,400 images of 90 building-scale indoor scenes. It provides surface reconstructions, globally aligned camera poses, and detailed 2D and 3D semantic segmentations at both region and object levels. The comprehensive panoramic coverage and precise global alignment enable a wide range of computer vision tasks, including keypoint matching, semantic segmentation, and view overlap prediction, particularly in the context of indoor scene understanding.
@@
Ego4D (Grauman et al., 2022): Ego4D is an unprecedented egocentric video dataset, comprising 3,670 hours of first-person footage collected by 931 participants across 74 locations in 9 countries. The dataset covers a wide array of daily-life scenarios and includes supplementary data such as audio, 3D meshes, gaze, and multi-camera views. With millions of annotations supporting various benchmark tasks, Ego4D enables research into first-person perception, episodic memory, and social interactions, significantly advancing the field of egocentric visual understanding.
@@
AgiBot-World (AgiBot-World-Contributors et al., 2025): AgiBot-World is a large-scale robotic manipulation dataset featuring over 1 million trajectories across 217 tasks in five real-world domains—including domestic, retail, industrial, restaurant, and office environments—collected by more than 100 real robots. Data collection is standardized and includes human-in-the-loop verification to ensure high quality and diversity. The dataset supports a wide range of hardware, from simple grippers to dexterous hands with visuo-tactile sensors, and provides rich multimodal annotations. AgiBot-World enables the development and evaluation of generalist robot policies in complex, long-horizon, and diverse scenarios, moving towards scalable, general-purpose robotic intelligence.
@@
DTU (Jensen et al., 2014): The DTU dataset is a large-scale multi-view stereo (MVS) benchmark designed for 3D reconstruction research. It contains 80 scenes with significant variability in surface properties and geometric complexity. Each scene is captured from 49 or 64 accurate camera positions and includes high-quality reference scans obtained via structured light. This dataset advances the scale and diversity compared to previous MVS datasets and serves as a challenging benchmark for evaluating the accuracy and completeness of stereo reconstruction algorithms.
@@
DAVIS 2017 (Pont-Tuset et al., 2017): DAVIS 2017 is a benchmark dataset for video object segmentation, consisting of 150 video sequences and a total of 10,459 annotated frames with pixel-level masks for multiple objects per scene. Compared to earlier versions, DAVIS 2017 introduces greater complexity, including multiple objects, occlusions, small and fine structures, and fast motion. The dataset has driven significant progress in video object segmentation and is widely used for benchmarking and competition in the field.
@@
Waymo Open Dataset (Sun et al., 2020): The Waymo Open Dataset is one of the largest and most diverse multimodal datasets for autonomous driving. It comprises 1,150 scenes, each lasting 20 seconds, captured across multiple cities with extensive geographic coverage. Each scene includes synchronized and calibrated high-resolution camera images and LiDAR data, along with exhaustive 2D and 3D bounding box annotations and consistent object tracking IDs. The dataset supports research in detection, tracking, and sensor fusion, providing a valuable foundation for developing robust and generalizable self-driving technologies.
MMSI-Bench is composed of questions from 8 data sources: 463 from Matterport3D, 280 from ScanNet, 67 from Ego4D, 45 from AgiBot-World, 72 from DTU, 39 from nuScenes, 29 from DAVIS 2017, and 5 from Waymo. We analyzed widely used VLM pre-training datasets, CC3M (Changpinyo et al., 2021) and LAION-COCO (Schuhmann et al., 2022), which are commonly utilized by open-source VLMs such as LLaVA, NVILA, and Qwen-VL. Our analysis indicates that outdoor scenes constitute only an estimated 10% of the data in these foundational VLM datasets. To align our benchmark with the typical training distribution of these models, we have curated MMSI-Bench to maintain a comparable proportion, with approximately 12% of our samples sourced from outdoor datasets. Recognizing that spatial understanding is one of the most critical capabilities for autonomous driving, we selected Waymo and nuScenes as the primary sources for these outdoor samples.
F.2 Data Annotation and Quality Control Interfaces
@@ F.3 General Guidelines
F.4 Data Format and Structure
@@
F.5 Quality Control and Validation
@@
F.6 Difficulty Annotation
We recruited new human evaluators to answer MMSI-Bench and recorded their per-question completion times. Each question was independently assigned to two evaluators. We normalized completion times across evaluators and used the mean normalized time as the difficulty indicator, with longer times indicating higher difficulty. The resulting distribution exhibited a clear three-tier structure. We therefore labeled each question as Easy, Medium, or Hard. Any question answered incorrectly by the evaluators was directly assigned to the Hard tier. The Easy tier contains 605 questions with an average human completion time of approximately 40 seconds. The Medium tier contains 262 questions with an average time of approximately 110 seconds. The Hard tier contains 133 questions, defined as those taking longer than approximately 110 seconds on average or those answered incorrectly.
Appendix G Representative MMSI-Bench Samples from Each Category (Full Version)